# Forecasting age distribution of deaths across countries: Life expectancy and annuity valuation

### v2

@@ Abstract
We investigate two transformations within the framework of compositional data analysis for forecasting the age distribution of death counts. Drawing on age-specific period life-table death counts from 24 countries in the Human Mortality Database, we assess and compare the point and interval forecast accuracy of the two transformations. Enhancing the forecast accuracy of period life-table death counts holds significant value for demographers, who rely on such forecasts to estimate survival probabilities and life expectancy, and for actuaries, who use them to price temporary annuities across various entry ages and maturities. While our primary focus is on temporary annuities, we also consider long-term contracts that, particularly at higher entry ages, approximate lifetime annuities.
Keywords: centered log-ratio transformation; cumulative distribution function transformation; life-table death counts; principal component analysis; single-premium temporary annuity
@@ 1 Introduction
Actuaries and demographers have long been engaged in mortality modeling and forecasting, dating back to the early 20th century, as a response to the financial challenges posed by rapid mortality improvements. These improvements affected the accuracy of annuity calculations and the sustainability of government pension systems (Pollard, 1987). Forecasting life tables for annuitants was a key topic at the 5th International Congress of Actuaries in Berlin in 1906 (Cramér & Wold, 1935). In the field of statistical demography, numerous researchers have proposed innovative approaches to modeling and forecasting mortality by age, sex, and cause, across time and space. Special emphasis has been placed on statistical time-series extrapolation models (see, e.g., Booth, 2006; Booth & Tickle, 2008; Cairns et al., 2008; Shang et al., 2011; Basellini et al., 2023, for comprehensive reviews). Within mortality modeling, at least three instruments are commonly used: age-specific mortality rates, survival probabilities, and the age distribution of death counts. Bergeron-Boucher et al. (2019) provide a thorough analysis of these mortality instruments, highlighting their respective advantages and limitations. All three instruments can be derived from one another using standard life-table relationships (see Preston et al., 2001, Chapter 3). Computationally, the conversion among these instruments can be achieved using the LifeTable function in the MortalityLaws package (Pascariu, 2025).
@@
Among the three mortality instruments, we focus on life-table death counts and model their evolution over time. These counts reveal a clear shifting pattern, with deaths progressively moving from younger to older ages. Between cohort and period life-table death counts, we choose to work with the latter, as period life tables are not only complete but also reflect the prevailing mortality conditions within a given time frame (see also Oeppen, 2008; Bergeron-Boucher et al., 2017). A time series of period life-table death counts enables the analysis of temporal changes in key longevity indicators, such as the modal age at death (see, e.g., Canudas-Romo, 2010), as well as measures of lifespan variability, including the standard deviation, interquartile range, and Gini coefficient (see, e.g., Vaupel et al., 2011; van Raalte & Caswell, 2013).
@@
In demography, Oeppen (2008) introduced the use of principal component analysis to model and forecast the age distribution of deaths within a compositional data analysis framework, treating age-specific life-table death counts ($d_{x}$) as compositional data. As with all compositional data, these counts are subject to inherent constraints, typically bounded between zero and a fixed upper limit. Consequently, the sample space of compositional data is a simplex, defined as
@@
where $\mathbb{S}$ denotes a simplex, which is a (D1)(D-1)-dimensional subset of real-valued space $R^{D-1}$, denotes vector transpose, and $c$ is a fixed constant, set typically to one (portions) (Scealy et al., 2015), $10^{6}$ parts per million (Scealy et al., 2015), or $10^{5}$ life-table death counts.
Due to the non-negativity and summability constraints inherent in compositional data, conventional linear techniques may be inappropriate, as they do not account for the unique variance-covariance structure of such data. A common strategy is to first transform the data to remove the summability constraint, thereby enabling the application of linear methods. One widely used approach is the family of log-ratio transformations, particularly the centered log-ratio (CLR) transformation (see, e.g., Aitchison, 1986). An alternative is the cumulative distribution function (CDF) transformation, which leverages monotonicity to effectively handle the presence of zero values (see Shang & Haberman, 2025b).
This paper makes two key contributions. First, we evaluate and compare the point forecast accuracy of the CLR and CDF transformations, using data from 24 countries in the Human Mortality Database (2025), with time series starting in or before 1950. Second, while the existing literature has largely focused on point forecasts, we implement a computationally efficient approach of Shang & Haberman (2025a) for constructing pointwise prediction intervals and compare the interval forecast accuracy between the CLR and CDF transformations. From a forecasting perspective, this study presents a comprehensive comparison across countries and further explores heterogeneity in mortality and its implications for life expectancy forecasting and annuity pricing.
The remainder of this paper is structured as follows. Section 2 describes the multi-country data sets used in the analysis. Section 3 revisits the CLR and CDF transformations and introduces a forecasting method based on principal component analysis. In Section 4, we assess point forecast accuracy using the Kullback-Leibler and Jensen-Shannon divergences. For evaluating interval forecast accuracy, we consider the empirical coverage probability and its absolute deviation from the nominal coverage level, as well as the interval score proposed by Gneiting & Raftery (2007). Section 5 applies the CDF transformation to estimate the single-premium temporary immediate annuity prices for various entry ages and maturities for female and male populations in 24 countries. Section 6 concludes with a summary and suggestions for future extensions of the proposed methodology.
@@ 2 Age distribution of death counts
The data sets used in this study are sourced from the Human Mortality Database (2025). We consider single-year ages ranging from 0 to 110. Out of 41 predominantly developed countries, we select 24 countries for which life-table death counts are available from 1950 or earlier. This results in 48 sex-specific populations used for analysis and comparison. The selected countries and their corresponding sample periods are listed in Table 1.
Table 1: Countries and their respective sample periods. For the annuity price calculations, we use the official interest rate as of 18 May 2025.
@@
The life-table death counts are non-negative and sum to a radix (i.e., a population experiencing 100,000 births annually) for each year. For the life-table death counts, there are 111 ages, and these are ages 0, 1, $\dots$, 109, 110+. Due to rounding, there may be zero counts for higher ages at some years. To rectify this problem, we work with the probability of dying (i.e., $q_{x}$) and the life-table radix to recalculate our estimated death counts (up to six decimal places). In doing so, we obtain more detailed death counts than the ones reported.
To understand the main features of the data, Figure 1 displays rainbow plots of the female and male age-specific life-table death counts in Australia in a single-year group from 1921 to 2021. The time ordering of the curves follows the colour order of a rainbow, where data from the distant past are shown in red while the more recent data are shown in purple (see Hyndman & Shang, 2010, for other examples). Both figures demonstrate a decreasing trend in infant death counts and a typical negatively skewed distribution, where the peaks gradually shift to higher ages for both sexes. This shift is a source of longevity risk, which creates a major issue for insurers and government pension funds, especially in the selling and risk management of annuity products (see Denuit et al., 2007, for a discussion). Moreover, the spread of the distribution indicates lifespan variability. A decrease in variability over time can be observed directly and can be measured, with interquartile range or Gini coefficient (see, e.g., Wilmoth & Horiuchi, 1999; van Raalte & Caswell, 2013; Debón et al., 2017). With the decrease in infant deaths and increase in mortality at older ages, the spread of the distribution has been reduced.
Figure 1: Rainbow plots of age-specific life-table death counts from 1921 to 2021 in a single-year group. The oldest years are shown in red, with the most recent years in violet. Curves are ordered chronologically according to the colours of the rainbow.
@@ 3 Transformation
Let the age-specific life-table death counts be denoted by $d^{s}_{t,i}$ or $d^{s}_{t,x}$, where $t$ denotes a year, $s$ represents females or males, and $i$ or $x$ denotes an age. For each year $t$, the life-table death counts sum to a radix $10^{5}$.
@@ 3.1 Centered log-ratio transformation
The CLR transformation presents a mapping from the simplex to a hyperplane in Euclidean space, where principal component analysis can be applied. The transformation consists of the following steps:
@@
    $\beta_{t,x}^{s}$ $=\ln\zeta^{s}_{t,x}$
    $=\ln d_{t,x}^{s}-\ln\alpha_{n,x}^{s}$
@@
    By applying eigendecomposition to $\bm{\beta}_{x}^{s}$, we can express
@@
    By conditioning on the estimated principal components and observed data, the $h$-step-ahead forecast of $\beta_{n+h,x}^{s}$ can be estimated by
@@
    where $\widehat{\gamma}_{n+h|n,\ell}$ denotes the $h$-step-ahead forecast of the principal component scores. These forecasts can be obtained via a univariate time-series forecasting method, such as the exponential smoothing method of Hyndman et al. (2008). The automatic algorithm of Hyndman et al. (2002) can be used to select the optimal exponential smoothing model among 15 possible candidates (see Table 1 Hyndman & Khandakar, 2008), based on the corrected Akaike information criterion.
@@
    Then, we add back the geometric mean to obtain the forecasts of the life-table death count $d_{n+h,x}^{s}$:
@@ 3.2 Cumulative distribution function transformation
For a time series of CDFs, it is important to scale the life-table death counts to sum to one. We describe the transformation in the following steps:
@@
    Compute the empirical cumulative distribution function via the cumulative sum,
@@
    where $D^{s}_{t,111}=1$.
@@
    Perform the logistic transformation,
@@
    With a set of unconstrained data $\bm{Z}^{s}_{x}=(Z^{s}_{1,x},\dots,Z^{s}_{n,x})^{\top}$, we implement a univariate functional time-series forecasting method. By computing the sample variance, we express a stochastic process $Z^{s}_{t,x}$ as
@@
    where $\psi^{s}_{k,x}$ denotes the $k$th principal component for age $x$ and sex $s$, $\eta^{s}_{t,k}=\langle Z^{s}_{t,x},\psi^{s}_{k,x}\rangle$ is the estimated principal component score at time $t$, $\varepsilon^{s}_{t,x}$ denotes the model residual function for sex $s$ in year $t$, and $K_{s}$ denotes the number of retained principal components. We implement an eigenvalue ratio criterion of Li et al. (2020), as well as $K_{s}=6$ recommended in Hyndman et al. (2013).
@@
    By conditioning on the estimated functional principal components $\bm{\Psi}^{s}_{x}=(\psi^{s}_{1,x},\dots,\psi^{s}_{K_{s},x})$ and observed data $\bm{Z}_{x}^{s}=(Z_{1,x}^{s},\dots,Z_{n,x}^{s})$, the $h$-step-ahead forecast of $Z_{n+h,x}^{s}$ can be obtained
@@
    $\widehat{d}^{s}_{n+h|n,x}=\Delta^{x}_{i=1}\widehat{D}^{s}_{n+h|n,i},$

### v3

@@ Abstract
In this paper, we provide a comprehensive cross-country validation study of compositional mortality modeling and forecasting methods. Thus, we consider two one-to-one transformations: the cumulative distribution function and the centered log-ratio transformation in compositional data analysis. Between the two transformations, the cumulative distribution function provides a scale-free way to visualize the gender gap and cross-country heterogeneity in the probability of dying by sex and country. Drawing on age-specific period life-table death counts from 24 countries in the Human Mortality Database (2025), we assess and compare the point and interval forecast accuracy of the two transformations, using the same forecasting method. Enhancing the forecast accuracy of period life-table death counts is of significant value to demographers, who rely on such forecasts to estimate survival probabilities and life expectancy, and to actuaries, who use them to price annuities across various entry ages and maturities.
Keywords: centered log-ratio transformation; cumulative distribution function transformation; life-table death counts; multi-country comparison; principal component analysis; single-premium temporary annuity
@@ 1 Introduction
Actuaries and demographers have long been engaged in mortality modeling and forecasting, dating back to the early 20th century, as a response to the financial challenges posed by rapid mortality improvements. These improvements affected the accuracy of annuity calculations and the sustainability of government pension systems (Pollard, 1987). Forecasting life tables for annuitants was a key topic at the 5th International Congress of Actuaries in Berlin in 1906 (Cramér and Wold, 1935). In the field of statistical demography, numerous researchers have proposed innovative approaches to modeling and forecasting mortality by age, sex, and cause, across time and space. Special emphasis has been placed on statistical time-series extrapolation models (see, e.g., Booth, 2006; Booth and Tickle, 2008; Cairns et al., 2008; Shang et al., 2011; Basellini et al., 2023, for comprehensive reviews). Within mortality modeling, at least three instruments are commonly used: age-specific mortality rates, survival probabilities, and age distribution of death counts. Bergeron-Boucher et al. (2019) provides a detailed analysis of these mortality instruments, highlighting their respective advantages and limitations. All three instruments can be derived from one another using standard life-table relationships (see Preston et al., 2001, Chapter 3). Computationally, the conversion among these instruments can be achieved using the LifeTable function in the MortalityLaws package (Pascariu, 2025).
Among the three mortality instruments, we focus on life-table death counts ($d_{x}$) and model their evolution over time and entire ages from $x=0$ to 110+. The $d_{x}$ is designed for a synthetic cohort (often normalized), not registered deaths, and that once one works on the $d_{x}$ scale, the mapping to actuarial quantities is straightforward. In particular,
$l_{x+1}=l_{x}-d_{x},\qquad q_{x}=\frac{d_{x}}{l_{x}},\qquad p_{x}=1-q_{x},$
so survival probabilities and annuity factor follow directly from the number of people alive $l_{x}$ sequence (if needed, one can obtain age-specific mortality $m_{x}$ from the probability of dying $q_{x}$ using standard approximation in the LifeTable function).
@@
Since it does not depend on exposure to risk, these synthetic death counts reveal a clear shifting pattern, with deaths progressively moving from younger to older ages. Between cohort and period life-table death counts, we choose to work with the latter, as period life tables are not only complete, but also reflect the prevailing mortality conditions within a given time frame (see also Oeppen, 2008; Bergeron-Boucher et al., 2017). A time series of period life-table death counts enables the analysis of temporal changes in key longevity indicators, such as the modal age at death (see, e.g., Canudas-Romo, 2010), as well as measures of lifespan variability, including standard deviation, interquartile range, and Gini coefficient (see, e.g., Vaupel et al., 2011; van Raalte and Caswell, 2013).
@@
In demography, Oeppen (2008) introduced the use of principal component analysis to model and forecast the age distribution of deaths within a compositional data analysis framework, treating age-specific life-table death counts as compositional data. As with all compositional data, these counts are subject to inherent constraints, typically bounded between zero and a fixed upper limit. Consequently, the sample space of compositional data is a simplex, defined as
@@
where $\mathbb{S}$ denotes a simplex, which is a (D1)(D-1)-dimensional subset of real-valued space $R^{D-1}$, denotes vector transpose and $c$ is a fixed constant, set typically to one (portions) (Scealy et al., 2015), $10^{6}$ parts per million (Scealy et al., 2015), or $10^{5}$ in the life-table death counts.
Due to the non-negativity and summability constraints inherent in compositional data, conventional linear techniques may be inappropriate, as they do not account for the unique variance-covariance structure of such data. A common strategy is first to transform the data to remove the summability constraint, thereby enabling the application of linear methods. One widely used approach is the family of log-ratio transformations, particularly the centered log-ratio (CLR) transformation (see, e.g., Aitchison, 1986). An alternative is the cumulative distribution function (CDF) transformation, which leverages monotonicity to effectively handle the presence of zero values (see Shang and Haberman, 2025).
This paper makes three key contributions. Based on the early work of Shang and Haberman (2026, 2025), we evaluate and compare the point forecast accuracy of the CLR and CDF transformations, using data from 24 countries in the Human Mortality Database (2025), with time series starting in or before 1950. Although the existing literature has largely focused on point forecasts, we implement a computationally efficient approach for constructing pointwise prediction intervals and compare the interval forecast accuracy between the CLR and CDF transformations. From a forecasting perspective, this study presents a comprehensive comparison between countries and further explores the heterogeneity of mortality, as well as its implications for forecasting life expectancy and pricing annuities. In addition, the CDF transformation provides a scale-free measure to visually compare the probability of dying between genders and across countries.
The remainder of this paper is structured as follows. Section 2 describes the multi-country data sets used in the analysis. Section 3 revisits the CLR and CDF transformations and introduces a forecasting method based on principal component analysis. In Section 4, we introduce new visualization plots, based on the CDF transformation, to compare the age distribution of deaths between sexes in the same country and across countries. In Section 5, we assess point forecast accuracy using the Kullback-Leibler and Jensen-Shannon divergences. For evaluating interval forecast accuracy, we consider the empirical coverage probability and its absolute deviation from the nominal coverage level, as well as the interval score proposed by Gneiting and Raftery (2007). Section 6 applies the CDF transformation to estimate the single-premium temporary immediate annuity prices for various entry ages and maturities for female and male populations in 24 countries. Section 7 concludes with a summary and suggestions for future extensions of the proposed methodology.
@@ 2 Age distribution of death counts
The data sets used in this study are sourced from the Human Mortality Database (2025). We consider single-year ages ranging from 0 to 110. Of the 41 predominantly developed countries, we select 24 for which life-table death counts are available from 1950 or earlier. This results in 48 sex-specific populations used for analysis and comparison. The selected countries and their corresponding sample periods are listed in Table 1.
Table 1: Countries and their respective sample periods. For the annuity price calculations, we use the official interest rate as of May 18, 2025.
@@
The life-table death counts are non-negative and sum to a radix (i.e., a population experiencing 100,000 births annually) for each year. For the life-table death counts, there are 111 ages: 0, 1, $\dots$, 109, 110+. Due to rounding, there may be zero counts for higher ages in some years. To rectify this problem, we work with the probability of dying (i.e., $q_{x}$) and the radix of the life-table to recalculate our estimated death counts (up to six decimal places). In doing so, we obtain more detailed death counts than the ones reported. Note that $q_{x}$ is the mortality instrument studied by Cairns et al. (2006).
To understand the main features of the data, Figure 1 displays rainbow plots of the age-specific life-table death counts for females and males in Australia, grouped by single year, from 1921 to 2021. The time order of the curves follows the color order of a rainbow, where data from the distant past are shown in red, while the most recent data are shown in purple (see Hyndman and Shang, 2010, for other examples). Both figures demonstrate a decreasing trend in infant death counts and a typical negative skew, with the peaks gradually shifting to higher ages for both sexes. This shift is a source of longevity risk, creating a major issue for insurers and government pension funds, especially in the sale and risk management of annuity products (see Denuit et al., 2007, for a discussion). Moreover, the spread of the distribution indicates lifespan variability. A decrease in variability over time can be directly observed and quantified using the interquartile range or the Gini coefficient (see, e.g., Wilmoth and Horiuchi, 1999; van Raalte and Caswell, 2013; Debón et al., 2017). With the decrease in infant deaths and the increase in mortality at older ages, the spread of the distribution has narrowed.
Figure 1: Rainbow plots of Australian age-specific life-table death counts from 1921 to 2021 in a single-year group. The oldest years are shown in red, with the most recent years in violet. Curves are ordered chronologically according to the colors of the rainbow.
While we present the Australian age distribution of deaths in Figure 1, in Section 4, we introduce new visualization plots for pairwise comparisons of the age distribution of death counts across sexes and countries.
@@ 3 Transformation
Let age-specific life-table death counts be denoted by $d^{s}_{t,i}$ or $d^{s}_{t,x}$, where $t$ denotes a year, $s$ represents females or males, and $i$ or $x$ denotes an age. For each year $t$, the life-table death counts sum to a radix $10^{5}$ or one by normalization.
@@ 3.1 Centered log-ratio transformation
The CLR transformation maps the simplex to a hyperplane in Euclidean space, allowing principal component analysis to be applied. The transformation consists of the following steps:
@@
    $\beta_{t,x}^{s}=\ln\zeta^{s}_{t,x}$ $=\ln d_{t,x}^{s}-\ln\alpha_{n,x}^{s}$
@@
    Applying eigendecomposition to the sample variance of $\bm{\beta}_{x}^{s}$, we can express
@@
    Forecasting step: By conditioning on the estimated principal components and observed data, the $h$-step-ahead forecast of $\beta_{n+h,x}^{s}$ can be estimated by
@@
    where $\widehat{\gamma}_{n+h|n,\ell}$ denotes the $h$-step-ahead forecast of the principal component scores. These forecasts can be obtained via a univariate time-series forecasting method, such as the exponential smoothing method of Hyndman et al. (2008). The automatic algorithm of Hyndman et al. (2002) can be used to select the optimal exponential smoothing model among 15 possible candidates (see Table 1 Hyndman and Khandakar, 2008), based on the corrected Akaike information criterion.
@@
    Then, we add back the geometric mean to obtain the $h$-step-ahead forecasts of the life-table death count $d_{n+h,x}^{s}$:
@@ 3.2 Cumulative distribution function transformation
For a time series of CDFs, it is essential to scale the life-table death counts so that they sum to one. We describe the transformation in the following steps:
@@
    Compute the empirical cumulative distribution function by the cumulative sum,
@@
    where $D^{s}_{t,111}=1$. Here, we relabel ages as $1,\dots,111$ to represent actual ages $0,\dots,109,110+$.
@@
    Perform the logit transformation of $D_{t,x}^{s}$ for all ages except the last one,
@@
    With a set of unconstrained data $\bm{Z}^{s}_{x}=(Z^{s}_{1,x},\dots,Z^{s}_{n,x})^{\top}$, we implement a univariate functional time-series forecasting method. By computing the sample variance of $\bm{Z}^{s}_{x}$, we express a stochastic process $Z^{s}_{t,x}$ as
@@
    where $\psi^{s}_{k,x}$ denotes the $k$th principal component for age $x$ and sex $s$, $\eta^{s}_{t,k}=\langle Z^{s}_{t,x},\psi^{s}_{k,x}\rangle$ is the estimated principal component score at time $t$, $\varepsilon^{s}_{t,x}$ denotes the model residual function for sex $s$ in year $t$, and $K_{s}$ denotes the number of principal components retained. We implement an eigenvalue ratio criterion of Li et al. (2020), as well as $K_{s}=6$ recommended in Hyndman et al. (2013).
@@
    Forecasting step: By conditioning on the estimated functional principal components $\bm{\Psi}^{s}_{x}=(\psi^{s}_{1,x},\dots,\psi^{s}_{K_{s},x})$ and observed data $\bm{Z}_{x}^{s}=(Z_{1,x}^{s},\dots,Z_{n,x}^{s})$, the $h$-step-ahead forecast of $Z_{n+h,x}^{s}$ can be obtained
@@
    $\widehat{d}^{s}_{n+h|n,x}=\left\{\begin{array}[]{ll}\widehat{D}^{s}_{n+h|n,1}&x=1\\ \Delta^{x}_{i=1}\widehat{D}^{s}_{n+h|n,i}=\widehat{D}^{s}_{n+h|n,x}-\widehat{D}^{s}_{n+h|n,x-1}&2\leq x\leq 111\end{array}\right$.
@@
In Sections 3.1 and 3.2, the forecasting steps are designed to model and forecast each series individually. This method does not account for the correlations between female and male data within the same country. Although explicitly modeling the correlation between multiple series can improve forecast accuracy, this is not the focus of this paper, as a similar study has been presented in Shang and Jiménez-Varón (2026).
4 Visualization plots of the gap in the age distribution of deaths
While the two transformations work with life-table death counts, the CDF transformation is advantageous for visualizing differences between sexes and countries. In the first step of the CDF transformation, the age distribution of deaths for a given year is converted into a CDF, providing a scale-free means of comparing different populations in terms of their probabilities of dying, commonly denoted $q_{x}$. By computing the differences, represented by
$G_{t}(u)=CDF^{\text{M}}_{t}(u)-\text{CDF}^{\text{F}}_{t}(u),$
between the two populations, we can visualize the gap in the probability of dying for Australia, shown in Figure 2. The gap between the Australian female and male populations is largest between 1950 and 2000, particularly among individuals aged 60 to 90. Since the lower bound of the gap is 0, it indicates that the male population has a higher probability of dying than the female population.
Figure 2: Image plot showing time stochastic dominance map between the probabilities of dying between Australian males and females from 1921 to 2021.
We can summarize domain strength at each time $t$ using an integral measure, such as
$S_{t}=\int_{u=0}^{110+}\left[\text{CDF}^{\text{M}}_{t}(u)-\text{CDF}^{\text{F}}_{t}(u)\right]du,$ (1)
where the integral can be approximated via the trapezoidal rule. When $S_{t}>0$, it implies that females tend to live longer. The integral measure is closely related to the difference between life expectancy at birth. Recall that the life expectancy at birth can be expressed as
$e_{0}=\int_{0}^{\infty}S(u)du$
where $S(u)$ is the survival function to the last age considered. Since the survival function is the complement of the CDF, it can be shown that
$S_{t}$ $=\int_{u=0}^{110+}\left[\text{CDF}^{\text{M}}_{t}(u)-\text{CDF}^{\text{F}}_{t}(u)\right]du$
$=\int^{110+}_{u=0}[S_{t}^{\text{F}}(u)-S_{t}^{\text{M}}(u)]du$
$=(e_{0}^{\text{F}}-e_{0}^{\text{M}})_{t}.$
In Figure 3, we present a time series plot of $S_{t}$ to show the gap-integral measure between Australian females and males, aggregated by age.
Figure 3: A time series plot of the integral measure of the probabilities of dying between the Australian males and females from 1921 to 2021.
While Figures 2 and 3 compare Australian females and males, the CDF, being a scale-free measure, can also be used to compare the probability of dying across countries. In Figure 4, we present image plots comparing the probability of dying of females and males between the UK and Australia. For comparison, UK females generally have a higher probability of dying than Australian females, except between 1950 and 1970, especially between ages 30 and 80. In addition, UK males have a higher probability of dying than Australian males, except between 1950 and 2000, especially between ages 20 and 60.
Figure 4: Image plot showing time stochastic dominance map between the probabilities of dying between the UK and Australian males and females from 1921 to 2021.
Using the integral measure in (1), we present two time series plots showing the gap integral measure between the UK and Australian females and males. Except for 1950 and 1975, the UK population generally has a higher probability of dying than the Australian population.
(a) Female data
(b) Male data
Figure 5: Time series plots of the integral measure of the probabilities of dying between the UK and Australian males and females from 1921 to 2021.